# Variety and Quality over Quantity: Towards Versatile Instruction Curation

### 12 Dec 2023 v1

@@ 4.1.1 Basic Foundation Models and Base Datasets
The instruction tuning implementation details for both tasks can be found in Appendix D
@@ 4.2 Benchmarks and Metrics
We have selected six widely-used benchmarks across the two tasks. For code generation tasks, we incorporate HumanEval and MBPP. And for NLU tasks, our selected benchmarks include HellaSwag, ARC Challenge, TruthfulQA, and MMLU. The details of these selected benchmarks can be found in Appendix E.1.
@@ 4.3 Experiment Results
We compare models fine-tuned on our paradigm’s final curated dataset with other state-of-the-art (SOTA) pretrained LLMs and instruction-tuned LLMs across both tasks. The details of the selected models for comparison is presented in Appendix E.2
@@ 4.3.1 Code Generation Tasks
As depicted in Table 1, our paradigm’s fine-tuned model outperforms most of the models in code generation tasks. While our model is nearly 2% lower than the current state-of-the-art 15B model, WizardCoder, on both benchmarks, it’s important to note that our paradigm utilizes only about one-eighth of the instruction data used by WizardCoder. Considering the size of the instruction dataset, our paradigm demonstrates robust performance, showcasing the ability to achieve performance close to the state-of-the-art with a significantly smaller amount of data.
@@
Model Fine-tuning Data Size HumanEval MBPP
CodeT5+ 16B Pretrained 0.309 -
Code LLaMA-13B 0.360 0.470
StarCoder-15B 0.336 0.436
@@
Code LLaMA Instruct-13B approx. 5B tokens 0.427 0.494
WizardCoder-15B 78k instructions 0.573 0.518
@@ 4.3.2 Natural Language Understanding Tasks
Ours-7B 15k instructions 0.844 0.643 0.490 0.645
Table 2 presents the NLU task comparison results. Notably, our final instruction-tuned model consistently outperforms other 7B models on all benchmarks. Even among 13B models, our model outperforms in all benchmarks except TruthfulQA. With only 7 billion parameters and 15k instructions, significantly fewer than other instruction-tuned models, our model achieves the highest average benchmark score at 0.656.
@@ 4.4 Paradigm Ablation Experiments Results
Our paradigm ablation experiment begins with the original base dataset serving as the input for our paradigm. Subsequently, we generate the expanded dataset, variety-compressed dataset, and the quality-compressed dataset. These datasets are then utilized for fine-tuning the basic foundation models. We assess the benchmark performance of these models to validate the effectiveness of each component of our paradigm.
@@ 4.4.1 Code Generation Tasks
Table 3 provides an overview of the paradigm experiments conducted on code generation tasks. For data expansion, we repeatedly perform the expansion step two times, resulting in a 60k size instruction dataset.
@@ 4.4.2 Natural Language Understanding Tasks
Table 4 presents our paradigm experiment results on four NLU benchmarks. For data expansion, we iteratively perform the expansion step three times, resulting in a 100k size instruction dataset.
@@ 5.1 Composition of The Final Curated Dataset
We take a step further to analyze the composition of the final curated dataset, unraveling the origins of diverse and high-quality instruction items. Figure 5 presents the source proportions of the final curated dataset for code generation and NLU tasks, yielding several noteworthy conclusions.
@@ 5.2 GPU Hours and Carbon Emission
By compressing the size of the instruction dataset, we aim to reduce the GPU hours required for instruction tuning, contributing to a subsequent decrease in carbon emissions. Table 5 illustrates the impact of different dataset sizes on GPU hours and CO2 emissions. GPU hours are calculated under the same settings of training epoch and batch size, while carbon emissions are computed using an online machine learning CO2 calculator [^1] https://mlco2.github.io/impact/#co2eq.
@@
5.3 Limitations
Redundancy in Data Expansion: The first step involves data expansion, introducing some data redundancy as the new instructions are primarily based on rewriting the original ones. Despite mitigation in the Variety Compress step, the impact of duplicated instructions from the expansion phase should be considered.
Subjectivity in Quality Evaluation: Quality evaluation relies heavily on the GPT Quality Evaluation, and despite careful criteria, additional statistical analysis beyond the length factor could enhance the precision of high-quality data selection.
Unexplored Compression Limits: Our study does not explicitly investigate the compression limits of the original dataset. Understanding compression limits for both datasets is essential for optimizing the trade-off between size and performance.
5.4 Future Works
Diverse Instruction Evolution Strategies: We aim to explore strategies for expanding data that are less reliant on the original dataset, utilizing techniques such as prompt engineering and implementing stringent data supervision strategies for generating diverse instructions.
Comprehensive Metrics for Quality Evaluation: Future work will involve enhancing quality evaluation metrics beyond the current focus on instruction length. This includes integrating more comprehensive metrics like information entropy analysis, coherence analysis, and word frequency analysis to provide a nuanced assessment of instruction quality.
In-Depth Exploration of Compression Limits: Further investigation into the compression limits of the original instruction dataset is also a focus. We will delve into understanding whether the observed robust performance extends to smaller curated datasets and explore the detailed relationship between compression ratios and instruction datasets.
@@ 6 Conclusions
This paper introduces a novel paradigm for curating large language model instruction datasets to enhance the quality and diversity of data for fine-tuning LLMs. The paradigm unfolds in three steps: Data Enhancement and Expansion, Variety Compress, and Quality Compress.
@@ Appendix C Quality Score Distribution
We have gathered the quality scores for the variety compressed dataset following the second step of our paradigm, both in code generation tasks and NLU tasks. The score distributions are depicted in Figure 6. Notably, the quality scores exhibit an approximately normal distribution within the score interval of 60 to 100 for both tasks. This observation validates the effectiveness of our scoring strategies in discerning low-quality data. It should be noted that the minor bumps near 0 stem from connection errors or OpenAI API calling ratio constraints, resulting in GPT scores of 0 for certain instructions.
@@ Appendix D Experiments Implementation Details
For both foundation models, we conduct training on Azure Machine Learning Studio’s cluster, utilizing 4 nodes, each equipped with 8 V100 GPUs featuring DeepSpeed Zero-3 (Rajbhandari et al., 2019) offload. Specifically, during the fine-tuning of Mistral 7B, we employ LoRA (Hu et al., 2021). This strategy is chosen for its ability to ensure a more steady convergence of loss, resulting in better performance. The detailed fine-tuning arguments are outlined in Table 6.
@@ E.2 Compared LLMs
  • Code LLaMA & Code LLaMA Instruct (Rozière et al., 2023). Code Llama is a code-specialized version of Llama2 (Touvron et al., 2023b) trained on code-specific datasets. Code LLaMA Instruct is a version further fine-tuned with human instructions and self-instruct code synthesis.

### 27 Dec 2023 v2

@@ 4.1.1 Basic Foundation Models and Base Datasets
The instruction tuning implementation details for both tasks can be found in Appx. D
@@ 4.2 Benchmarks and Metrics
We have selected six widely-used benchmarks across the two tasks. For code generation tasks, we incorporate HumanEval and MBPP. And for NLU tasks, our selected benchmarks include HellaSwag, ARC Challenge, TruthfulQA, and MMLU. The details of these selected benchmarks can be found in Appx. E.1.
@@ 4.3 Experiment Results
We compare models fine-tuned on our paradigm LIFT’s final curated dataset with other state-of-the-art (SOTA) pretrained LLMs and instruction-tuned LLMs across both tasks. The details of the selected models for comparison is presented in Appx. E.2
@@ 4.3.1 Code Generation Tasks
As depicted in Tab. 1, our paradigm’s fine-tuned model outperforms most of the models in code generation tasks. While our model is nearly 2% lower than the current state-of-the-art 15B model, WizardCoder, on both benchmarks, it’s important to note that our paradigm utilizes only about one-eighth of the instruction data used by WizardCoder. Considering the size of the instruction dataset, our paradigm demonstrates robust performance, showcasing the ability to achieve performance close to the state-of-the-art with a significantly smaller amount of data.
@@
Model Data Size HumanEval MBPP
CodeT5+ -* 0.309 -
CodeLLaMA 0.360 0.470
StarCoder 0.336 0.436
@@
WizardCoder 78k 0.573 0.518
@@ 4.3.2 Natural Language Understanding Tasks
LIFT-7B 15k instructions 0.844 0.643 0.490 0.645
Tab. 2 presents the NLU task comparison results. Notably, our final instruction-tuned model consistently outperforms other 7B models on all benchmarks. Even among 13B models, our model outperforms in all benchmarks except TruthfulQA. With only 7 billion parameters and 15k instructions, significantly fewer than other instruction-tuned models, our model achieves the highest average benchmark score at 0.656.
@@ 4.4 Paradigm Ablation Experiments Results
Our paradigm ablation experiment begins with the original base dataset serving as the input for LIFT. Subsequently, we generate the expanded dataset, variety-compressed dataset, and the quality-compressed dataset. These datasets are then utilized for fine-tuning the basic foundation models. We assess the benchmark performance of these models to validate the effectiveness of each component of our paradigm.
@@ 4.4.1 Code Generation Tasks
Tab. 3 provides an overview of the paradigm experiments conducted on code generation tasks. For data expansion, we repeatedly perform the expansion step two times, resulting in a 60k size instruction dataset.
@@ 4.4.2 Natural Language Understanding Tasks
Tab. 4 presents our paradigm experiment results on four NLU benchmarks. For data expansion, we iteratively perform the expansion step three times, resulting in a 100k size instruction dataset.
@@ 5.1 Composition of The Final Curated Dataset
We take a step further to analyze the composition of the final curated dataset, unraveling the origins of diverse and high-quality instruction items. Fig. 4 presents the source proportions of the final curated dataset for code generation and NLU tasks, yielding several noteworthy conclusions.
@@ 5.2 GPU Hours and Carbon Emission
By compressing the size of the instruction dataset, we aim to reduce the GPU hours required for instruction tuning, contributing to a subsequent decrease in carbon emissions. Tab. 5 illustrates the impact of different dataset sizes on GPU hours and CO2 emissions. GPU hours are calculated under the same settings of training epoch and batch size, while carbon emissions are computed using an online machine learning CO2 calculator [^1] https://mlco2.github.io/impact/#co2eq.
@@
5.3 Limitations and Future Works
Our paradigm has limitations requiring further research for resolution.
Redundancy in Data Expansion: The first step involves data expansion, introducing some data redundancy as the new instructions are primarily based on rewriting the original ones. Despite mitigation in the Variety Compress step, the impact of duplicated instructions from the expansion phase should be considered. We aim to explore strategies for expanding data that are less reliant on the original dataset, implementing stringent data supervision strategies for generating diverse instructions.
Subjectivity in Quality Evaluation: Quality evaluation relies heavily on the GPT Quality Evaluation, and despite careful criteria, additional statistical analysis beyond the length factor could enhance the precision of high-quality data selection. Future work will involve enhancing quality evaluation metrics beyond the current focus on instruction length.
@@ 6 Conclusions
This paper introduces a novel paradigm LIFT for curating large language model instruction datasets to enhance the quality and diversity of data for fine-tuning LLMs. The paradigm unfolds in three steps: Data Enhancement and Expansion, Variety Compress, and Quality Compress.
@@ Appendix C Quality Score Distribution
We have gathered the quality scores for the variety compressed dataset following the second step of our paradigm, both in code generation tasks and NLU tasks. The score distributions are depicted in Fig. 5. Notably, the quality scores exhibit an approximately normal distribution within the score interval of 60 to 100 for both tasks. This observation validates the effectiveness of our scoring strategies in discerning low-quality data. It should be noted that the minor bumps near 0 stem from connection errors or OpenAI API calling ratio constraints, resulting in GPT scores of 0 for certain instructions.
@@ Appendix D Experiments Implementation Details
For both foundation models, we conduct training on Azure Machine Learning Studio’s cluster, utilizing 4 nodes, each equipped with 8 V100 GPUs featuring DeepSpeed Zero-3 (Rajbhandari et al., 2019) offload. Specifically, during the fine-tuning of Mistral 7B, we employ LoRA (Hu et al., 2021). This strategy is chosen for its ability to ensure a more steady convergence of loss, resulting in better performance. The detailed fine-tuning arguments are outlined in Tab. 6.
@@ E.2 Compared LLMs
  • Code LLaMA (Rozière et al., 2023). Code Llama is a code-specialized version of Llama2 (Touvron et al., 2023b) trained on code-specific datasets.